OpenStudy (anonymous):

Find the roots of x^3-2x^2+10x+136=0

OpenStudy (anonymous):

\[x ^{3}-2x ^{2}+10x+136=0\]

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ (x+4)(x^2-6x+34)=0\LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

I factored it, can you do it from here ?

OpenStudy (dangerousjesse):

Solve for x: \[x^3-2 x^2+10 x+136 = 0 \] Factor the left hand side. The left hand side factors into a product with two terms: \[(x+4) (x^2-6 x+34) = 0 \] Solve each term in the product separately. Split into two equations: \[x+4 = 0 \] or \[x^2-6 x+34 = 0 \] Look at the first equation: Solve for x. Subtract 4 from both sides: \[x = -4 \] or \[x^2-6 x+34 = 0 \] Look at the second equation: Solve the quadratic equation by completing the square. Subtract 34 from both sides: \[x = -4 \] or \[x^2-6 x = -34 \] Take one half of the coefficient of x and square it, then add it to both sides. Add 9 to both sides: \[x = -4 \] or \[x^2-6 x+9 = -25 \] Factor the left hand side. Write the left hand side as a square: \[x = -4 \] or \[(x-3)^2 = -25 \] Eliminate the exponent on the left hand side. Take the square root of both sides: \[x = -4 \] or \[x-3 = 5 i \] or \[x-3 = -5 i \] Look at the second equation: Solve for x. Add 3 to both sides: \[x = -4 \] or \[x = 3+5 i \] or \[x-3 = -5 i \] Look at the third equation: Solve for x. Add 3 to both sides:

OpenStudy (solomonzelman):

99% of the work:)

OpenStudy (solomonzelman):

I got suspended for this number of times

OpenStudy (dangerousjesse):

Pardon?

OpenStudy (solomonzelman):

I don't mind

OpenStudy (solomonzelman):

I mean, \(\LARGE\color{black}{ \rm I\LARGE\color{white}{ \rm │ }}\)don't mind.

OpenStudy (solomonzelman):

I am also a fan of completing the square, btw.

OpenStudy (anonymous):

Thanks :D

OpenStudy (dangerousjesse):

No problem, what'd you get?

OpenStudy (anonymous):

x=-5i+3, x=5i+3, x=-4

OpenStudy (anonymous):

Or you could do x=3+5i, x=3-5i for the first 2.